OpenStudy (anonymous):

a beryllium nucleus(mass 1.51*10^-26kg) is moving with a speed of 3.0*10^6m/s when it collides a elastically with a stationary deuterium nucleus of mass 3.34*10^-27kg. After the collision both nuclei moves off in the same direction.Determine the velocity of each nucleus after the collision.

OpenStudy (anonymous):

since there are no external forces whe we consider the neuclei and deuterium as a system, we can apply law of conservation of linear momentum (1.51*10^-26kg)*(3.0*10^6m/s)=(15.1*10^-27kg + 3.34*10^-27kg) * v (4.53*10^-20)=(18.44*10^-20)*v v = (4.53*10^-20)/ (18.44*10^-20) =(4.53/18.44)*10^7 =0.2456*10^7 24.56*10^5 m/s